three squares have sides with lengths a,b and c. if b is 20%greater than a, c is 25%greater than b, then by what percent is the area of the largest square greater than the area of the smallest square
Answer by sachi(548)

three squares have sides with lengths a,b and c.
if b is 20%greater than a, c is 25%greater than b,
so b=a+20%a=1.2a
& c=b+25%b=1.25b=1.25*1.2a=1.5a
so the sides of the 3 squares are a,1.2a & 1.5a
now it is clear that greatest square is of sides 1.5a & smallest is of a
so the areas are (1.5a)^2=2.25a^2 & a^2
the percent difference in the area of the largest square & the smallest square
=(2.25a^2-a^2)*100/a^2=(2.25-1)*100=125%